### v1

@@ 0.A.3.2 CelebA
As we have discussed in the main paper we employ a two phase training which consists of a pretraining stage and a counterfactual fine-tuning stage. We pretrain the model for 2M iterations with a batch size of 32 using the AdamW optimizer with initial learning rate $10^{-3}$, $\beta_{1}$=0.9, $\beta_{2}$=0.9, weight decay 0.01, and early stopping patience set to 10. We use a linear warmup scheduler for the first 100 iterations and also set values for gradient clipping and gradient skipping to 350 and 500 accordingly. We further fine-tune the model for 2K iterations with a initial learning rate of $10^{-4}$ and the same batch size and warmup scheduler, while we use a distinct AdamW optimizer with learning rate 0.1 to perform gradient ascent on the Lagrange coefficient.
0.A.4 GAN
@@
We adopt the experimental setup described in [3]. To condition the model, we directly pass the parent variables into the networks by concatenating them with the image for the encoder and discriminator, and with the latent variable for the generator.
0.A.4.1 MorphoMNIST
@@
In Tables 11, 12 and 13 we show the details of encoder, generator and discriminator architectures, respectively, for generating MorphoMNIST counterfactuals. In order to monitor the generative abilities of the model during training, we compare the factual images $x$ with the generated samples produced from a random noise latent $z$ and the ones produced from the latent representations of the encoder. We do this by measuring the $l_{1}$ distance between their embeddings, which we get by concatenating the last layer activations across all our predictors. For monitoring the fine-tuning of the GAN encoder, we measure the $l_{1}$ distance between the factual image embeddings and the ones of the samples generated by the encoder-produced latents.
Table 11: GAN Encoder Architecture for MorphoMNIST.
Layer F K S P A
Conv2D 64 4 2 1 LReLU(0.2)
Conv2D 128 4 2 1 LReLU(0.2)
Conv2D 256 4 2 1 LReLU(0.2)
Conv2D 512 1 2 1 LReLU(0.2)
Conv2D 512 1 2 1 None
Table 12: GAN Generator Architecture for MorphoMNIST.
Layer F K S P A
Transposed Conv2D 256 3 1 0 LReLU(0.2)
Transposed Conv2D 512 4 2 0 LReLU(0.2)
Transposed Conv2D 256 3 2 1 LReLU(0.2)
Transposed Conv2D 128 3 2 1 LReLU(0.2)
Transposed Conv2D 64 4 1 0 Tanh
Table 13: GAN Discriminator Architecture for MorphoMNIST.
Layer F K S P BN D A
$D_{z}$
Conv2D 512 1 1 0 N 0.2 LReLU(0.1)
Conv2D 512 1 1 0 N 0.5 LReLU(0.1)
$D_{x}$
Conv2D 32 3 2 0 Y 0.2 LReLU(0.1)
Conv2D 64 3 2 0 Y 0.2 LReLU(0.1)
Conv2D 128 3 2 0 Y 0.5 LReLU(0.1)
Conv2D 256 3 2 0 Y 0.5 LReLU(0.1)
Conv2D 512 3 2 0 N 0.5 LReLU(0.1)
$D_{xz}$
Conv2D 1024 4 1 0 N 0.2 LReLU(0.1)
Conv2D 1024 4 1 0 N 0.2 LReLU(0.1)
Conv2D 1 4 1 0 N 0.2 Sigmoid
For both training and fine-tuning, we use the Adam optimizer with initial learning rate $10^{-4}$, $\beta_{1}$=0.5, $\beta_{2}$=0.999 and batch size of 128. We implement early stopping with a patience of 10 epochs.
0.A.4.2 CelebA
@@
The architectures of encoder, generator and discriminator for CelebA are presented in Tables 14, 15 and 16, respectively. We monitor the generative abilities of the model during training by measuring the FID score of the samples generated from random noise latent variables $z$ and latents produced by the encoder.
Table 14: GAN Encoder Architecture for CelebA.
Layer F K S P BN A
Conv2D 64 2 1 0 Y LReLU(0.2)
Conv2D 128 7 2 0 Y LReLU(0.2)
Conv2D 256 5 2 0 Y LReLU(0.2)
Conv2D 256 7 2 0 Y LReLU(0.2)
Conv2D 512 4 1 0 Y LReLU(0.2)
Conv2D 512 1 1 0 Y LReLU(0.2)
Conv2D 512 1 1 0 Y None
Table 15: GAN Generator Architecture for CelebA.
Layer F K S P BN A
Transposed Conv2D 512 4 1 0 Y LReLU(0.2)
Transposed Conv2D 256 7 2 0 Y LReLU(0.2)
Transposed Conv2D 256 5 2 0 Y LReLU(0.2)
Transposed Conv2D 128 7 2 0 Y LReLU(0.2)
Transposed Conv2D 64 2 1 0 Y LReLU(0.2)
Transposed Conv2D 3 1 1 0 Y Sigmoid
Table 16: GAN Discriminator Architecture for CelebA.
Layer F K S P BN D A
$D_{z}$
Conv2D 1024 1 1 0 N 0.2 LReLU(0.1)
Conv2D 1024 1 1 0 N 0.5 LReLU(0.1)
$D_{x}$
Conv2D 64 3 2 0 Y 0.2 LReLU(0.1)
Conv2D 128 3 2 0 Y 0.2 LReLU(0.1)
Conv2D 256 3 2 0 Y 0.5 LReLU(0.1)
Conv2D 256 3 2 0 Y 0.5 LReLU(0.1)
Conv2D 1024 3 2 0 N 0.5 LReLU(0.1)
$D_{xz}$
Conv2D 2048 4 1 0 N 0.2 LReLU(0.1)
Conv2D 2048 4 1 0 N 0.2 LReLU(0.1)
Conv2D 1 4 1 0 N 0.2 Sigmoid
@@
As in the experiments on MorphoMNIST, we monitor fine-tuning in CelebA using the LPIPS metric (on VGG-16) between real images and samples generated by the encoder-produced latents. The rest training parameters adhere to the same configuration as described for MorphoMNIST.
0.A.5 Anti-causal predictors
0.A.5.1 MorphoMNIST
We use the implementation of [4], which is displayed in Table 17. We train deep convolutional regressors for thickness and intensity respectively, and a deep convolutional classifier for the digit variable. Specifically, as thickness affects both the image and the intensity, the anti-causal predictor of thickness should be conditioned on intensity. We can denote the thickness predictor as $q_{\theta}(t|x,i)$. Intensity and digit affect only the image, hence the distributions that the corresponding predictors model can be expressed as $q_{\theta}(i|x)$, $q_{\theta}(y|x)$. Both the regressors and the classifier are trained for 10K iterations with batch size 256 using the AdamW optimizer with initial learning rate $10^{-3}$ and a linear warmup scheduler for the first 100 iterations.
Table 17: Predictor architecture for MorphoMNIST. num_outputs represents the dimension of the output vector (1 for the regressors, 10 for the classifier).
Layer F K S P BN A
Conv2D 16 7 1 3 Y LReLU(0.1)
Conv2D 32 3 2 1 Y LReLU(0.01)
Conv2D 32 3 1 1 Y LReLU(0.01)
Conv2D 64 3 2 1 Y LReLU(0.01)
Conv2D 64 3 1 1 Y LReLU(0.01)
Conv2D 128 3 2 1 Y LReLU(0.01)
Linear 512 - - - Y LReLU(0.01)
Linear num_outputs - - - N -
0.A.5.2 CelebA
For each of the attributes smiling and eyeglasses we use the architecture of the binary convolutional classifier as defined in [30]. The architecture is also depicted in Table 18. We train each classifier $q_{\psi}(s|x)$, $q_{\psi}(e|x)$ for 25K iterations with a batch size of 128 leveraging the AdamW optimizer with a initial learning rate of $10^{-3}$ and a linear warmup scheduler. Additionally, for data augmentation we perform random horizontal flipping, and random cropping. Finally, similarly to [30], we choose to use a weighted random sampler (with replacement) to balance the distribution of the classes in each training batch.
Table 18: Classifier architecture for CelebA.
Layer F K S P BN A
Conv2D 16 3 1 1 Y ReLU
Conv2D 32 3 2 1 Y ReLU
Conv2D 32 3 1 1 Y ReLU
Conv2D 64 3 2 1 Y ReLU
Conv2D 64 3 1 1 Y ReLU
Conv2D 128 3 2 1 Y ReLU
AvgPool2D - 1 - - - -
Linear 128 - - - Y ReLU
Linear 1 - - - N Sigmoid
Appendix 0.B Qualitative results
0.B.1 MorphoMNIST
(a) VAE ($\sigma=\sqrt{e^{-5}}$)
(b) VAE ($\sigma$ learned)
(c) HVAE
(d) GAN
(e) GAN fine-tuned
Figure 6: Composition for 10 cycles on the MorphoMNIST test set.
(a) VAE($\sigma=\sqrt{e^{-5}}$)
(b) VAE ($\sigma$ learned)
(c) HVAE
(d) GAN
(e) GAN fine-tuned
Figure 7: Interventions on MorphoMNIST.
0.B.2 CelebA
(a) VAE ($\sigma=\sqrt{e^{-5}}$)
(b) VAE ($\sigma$ learned)
(c) HVAE fine-tuned
(d) GAN
(e) GAN fine-tuned
Figure 8: Composition for 10 cycles on the CelebA test set.
(a) VAE($\sigma=\sqrt{e^{-5}}$)
(b) VAE($\sigma$ learned)
(c) HVAE fine-tuned
(d) GAN
(e) GAN fine-tuned
Figure 9: Interventions on CelebA.

### v5

@@ B.4 HVAE
As we have discussed in the main paper we employ a two phase training which consists of a pretraining stage and a counterfactual fine-tuning stage. We pretrain the model for 2M iterations with a batch size of 32 using the AdamW optimizer with initial learning rate $10^{-3}$, $\beta_{1}$=0.9, $\beta_{2}$=0.9, weight decay 0.01, and early stopping patience set to 10. We use a linear warmup scheduler for the first 100 iterations and also set values for gradient clipping and gradient skipping to 350 and 500 accordingly. We further fine-tune the model for 2K iterations with a initial learning rate of $10^{-4}$ and the same batch size and warmup scheduler, while we use a distinct AdamW optimizer with learning rate 0.01 to perform gradient ascent on the Lagrange coefficient.
For ADNI, we utilised the architecture used in (7) for another dataset of brain MRIs (UK Biobank (49)). We train the model with the same hyperparameters detailed above for CelebA. We find that the counterfactual fine-tuning step is not needed since the model can produce counterfactuals and does not learn to just reproduce the original image. We tried adding this step, but found it to harm the performance of the model.
B.5 GAN
@@
For MorphoMNIST and CelebA experiments, we adopt the experimental setup described in (6). To condition the model, we directly pass the parent variables into the networks by concatenating them with the image for the encoder and discriminator, and with the latent variable for the generator. For both training and fine-tuning, we use the Adam optimizer with initial learning rate $10^{-4}$, $\beta_{1}$=0.5, $\beta_{2}$=0.999 and batch size of 128. We implement early stopping with a patience of 10 epochs.
@@
In order to monitor the generative abilities of the model for the MorphoMNIST dataset, during training, we compare the factual images $x$ with the generated samples produced from a random noise latent $z$ and the ones produced from the latent representations of the encoder. We do this by measuring the $l_{1}$ distance between their embeddings, which we get by concatenating the last layer activations across all our predictors. For monitoring the fine-tuning of the GAN encoder, we measure the $l_{1}$ distance between the factual image embeddings and the ones of the samples generated by the encoder-produced latents.
@@
For CelebA, we monitor the generative abilities of the model during training by measuring the FID score of the samples generated from random noise latent variables $z$ and latents produced by the encoder. As in the experiments on MorphoMNIST, we monitor fine-tuning in CelebA using the LPIPS metric (on VGG-16) between real images and samples generated by the encoder-produced latents.
For ADNI, we initially tried adding more and wider layers to the architecture used for CelebA but the model would not converge even with extensive tuning of the hyperparameters. We modified the architecture of (59), which was used to generate brain MRIs, intervening on the age in the ADNI dataset. We should note that the original architecture and training procedure would learn to reproduce the image and therefore did not work with a complex SCM as in our case. We removed the residual connections of the generator and modified the conditioning mechanism to use a concatenation of three variables (encoded with Fourier embeddings as in (3)). Finally, we used the losses we discuss for Conditional GAN above to make for a fairer comparison across datasets.
@@
The rest training parameters for CelebA and ADNI experiments adhere to the same configuration as described for MorphoMNIST.
B.6 Anti-causal predictors
For the MorphoMNIST attributes, we train deep convolutional regressors for continuous variables and a deep convolutional classifier for the digit, using the implementation of (7). For CelebA (simple/complex) and ADNI attributes we used ResNet-18 pretrained on ImageNet. Additionally, as CelebA is a highly unbalanced dataset, similarly to (35) we choose to use a weighted sampler to ensure a balanced class distribution in each training batch. Moreover, the anti-causal predictor of parent variables, are conditioned on all their children variables, including the image. All predictors were trained with batch size 256 using the AdamW optimizer with initial learning rate $10^{-3}$ and a linear warmup scheduler for the first 100 iterations.
Appendix C Additional qualitative results
C.1 Composition
Figure 6: Composition for all datasets (10 cycles)
C.2 Effectiveness
Figure 7: Effectiveness qualitative results for all datasets
Figure 8: Celeba-HQ (complex) (resolution 256x256) with HVAE. Counterfactual training produces unrealistic artifacts on higher-resolutions.
Appendix D Additional quantitative results
Table 6: Effectiveness on ADNI test set for all variables.
F1 $\uparrow$ MSE $\downarrow$ F1 $\uparrow$ MSE $\downarrow$ MSE $\downarrow$ F1 $\uparrow$
Model $do(apoE)$ $do(age)$ $do(sex)$ $do(brain_{vol})$ $do(vent_{vol})$ $do(slice)$
VAE $0.278_{0.19}$ $0.169_{0.04}$ $0.589_{0.35}$ $0.167_{0.03}$ $0.199_{0.04}$ $0.408_{0.10}$
HVAE $0.280_{0.19}$ $0.170_{0.03}$ $0.589_{0.35}$ $\mathbf{0.096_{0.03}}$ $\mathbf{0.037_{0.01}}$ $\mathbf{0.409_{0.13}}$
GAN $0.291_{0.237}$ $0.151_{0.037}$ $0.544_{0.365}$ $0.173_{0.04}$ $0.223_{0.036}$ $0.045_{0.044}$
Table 7: F1 Scores of classifers on CelebA validation set.
Attribute (F1) $\uparrow$ ResNet-18 Standard CNN
Age 0.921 0.901
Gender 0.987 0.943
Beard 0.972 0.932
Bald 0.721 0.642
Appendix E Diffusion models
Here we describe our efforts to extend an existing method that performs counterfactual image generation using diffusion models, but does so for causal graphs with a single variable. We adapted the classifier-free guidance setting of (43) in order to model the causal graphs of the datasets we are benchmarking. Initially this method under-performed (which we confirmed with the authors as a limitation of their model) when modeling non-trivial graphs (with multiple variables), which can be partially attributed to the spurious correlation of the variables. Also, we empirically observed that the nature of the conditioning mechanisms used and the architectural and hyperparameter choices utilised, can make an important difference for the performance of diffusion models.
We incorporated a UNet2DConditionModel taken from the diffusers library (40) in our codebase and tested it using our framework. We followed the same setting as in all our models, treating the forward diffusion process as the noise abduction and the backward process as the prediction.
We include experiments on MorphoMNIST, as well as CelebA (simple). We report that for MorphoMNIST, the Diffusion model performs worse than the rest of the models in terms of composition and realism (FID), while its performance on effectiveness and minimality (CLD) is comparable to that of the GAN. This is somehow expected, as this dataset is not complex in terms of image content and our simpler models, such as VAE perform well. For CelebA (using the simple causal graph), the diffusion model performs excellent on composition for 1 cycle, but it deteriorates significantly after more cycles. Regarding effectiveness, its performance is slightly worse than the other models, but ranks second when measuring its effect of intervening on eyeglasses. Finally, the realism and minimality of generated images is fairly good but worse than the HVAE.
We believe that the underperformance of the current Diffusion model in the task we are benchmarking is related to architectural and hyperparameter choices utilised. One of them can be the cross-attention conditioning mechanism we incorporated (between the attributes and the image). As this extension is a naive approach, we believe that novel methodological contribution is needed to enable a more fair comparison with other generative models that have been already proposed for counterfactual image generation in this setting.
Table 8: Composition of Diffusion model with cross-attention on MorphoMNIST and CelebA (simple).
Model MorphoMNIST ($L_{1}$ image space) CelebA (LPIPS)
1 cycle 10 cycles 1 cycle 10 cycles
Diffusion 17.9783.310 18.9633.931 0.0810.010 0.5420.014
Table 9: Effectiveness of Diffusion model with cross-attention on MorphoMNIST and CelebA (simple).
MorphoMNIST
Model Thickness (t) MAE $\downarrow$ Intensity (i) MAE $\downarrow$ Digit (y) Acc. $\uparrow$
do(t) do(i) do(y) do(t) do(i) do(y) do(t) do(i) do(y)
Diffusion 0.1830.005 0.2810.01 0.1630.005 12.598 1.033 19.267 1.009 13.859 0.651 0.848 0.015 0.791 0.015 0.563 0.01
CelebA (simple)
Smiling (s) F1 $\uparrow$ Eyeglasses (e) F1 $\uparrow$
$do(s)$ $do(e)$ $do(s)$ $do(e)$
Diffusion 0.823 0.086 0.738 0.153 0.863 0.008 0.910 0.036
Table 10: Realism (FID) and Minimality (CLD) of Diffusion model with cross-attention on MorphoMNIST and CelebA (simple).
MorphoMNIST CelebA (simple)
Model FID $\downarrow$ CLD $\downarrow$ FID $\downarrow$ CLD $\downarrow$
Diffusion 60.349 0.282 29.620 0.299
Figure 9: Composition (10 cycles) for CelebA (simple) with Diffusion Model (conditioning with cross-attention)
Figure 10: Effectiveness for MorphoMNIST and CelebA (simple) with Diffusion Model (conditioning with cross-attention)